OpenStudy (anonymous):

Find standard form of 3y^2+6y+108x-969=0 The form should be (y-k)^2=4p(x-h)

OpenStudy (mertsj):

\[3(y^2+2y+1)=-108x+969+3\] \[3(y+1)^2=-108x+972\] \[(y+1)^2=-36x+324\] \[(y+1)^2=-36(x-9)\]